OpenStudy (narissa):

evaluate -3x divided by y

OpenStudy (narissa):

a)x=10 b)x=-10 y=3 c)x=-10 y=-3

OpenStudy (narissa):

for a )y =-3

OpenStudy (astrophysics):

Huh? \[-\frac{ 3x }{ y }\]

OpenStudy (astrophysics):

Could you explain what exactly you have to do? Are we plugging things in or what

OpenStudy (narissa):

yes

OpenStudy (narissa):

the expression is -3x divided by y and all three a b c i posted what has to be plugged in to it

OpenStudy (narissa):

so first i think looks like -3(10 divided by-3)=

OpenStudy (astrophysics):

Ok, so where are you having trouble, just plug in the values so you have x = 10, and y = -3, so we have \[-\frac{ 3(10) }{ (-3) } = \frac{ 3(10) }{ 3 } = 10\]

OpenStudy (astrophysics):

Do the same with the rest

OpenStudy (narissa):

idk how to do it

OpenStudy (narissa):

k thanks

OpenStudy (narissa):

@jango_IN_DTOWN

OpenStudy (astrophysics):

What do you mean? It's basic arithmetic

OpenStudy (narissa):

so it equals 10

OpenStudy (astrophysics):

Yes

OpenStudy (astrophysics):

Just write it out, and do the math :P

OpenStudy (narissa):

ok thanks

OpenStudy (narissa):

so b is -10

OpenStudy (narissa):

@Astrophysics